A spring mass damper system is fixed at one end. The damper behaves such that a constant force of 66 N applied to the damper gives a velocity of 7.95 m/second. Determine the damping constant 'c' ?
Answer #1

Damping force F = C * V

where C = Damping constant

& V = Velocity

so,

F = 66 N

V = 7.95 m/s

Therefore

66 = C * 7.95

=> C = 66/7.95

C = 8.30 N-s/m